OpenStudy (anonymous):

What is the slope of the line that contains the points (-7, 5) and (-2, 3)?

OpenStudy (anonymous):

This is a classic example of the "point-slope formula" from Algebra! \[y_2 – y_1 = m(x_2 – x_1)\] What you're doing is taking the second point's coordinates (x,y) and putting them in for the x2's and then subtracting by the initial point's x1's. See if you get slope (that's what 'm' stands for) is -2/5 or -0.4, ok?

OpenStudy (anonymous):

You line should be y=(-2/5)x+(11/5). Steps: 1. Input coordinates into x1, x2, y1, x2; in the order of the change (second minus first) 2. Solve for 'm' (slope) 3. Input 'm' back in but into the y=mx+b equation instead, and use either one of your two coordinates for the x & y values. 4. Solve for 'b' (y-intercept) 5. Rewrite with a generic x & y, using the known values now for m & b